# Graph-Informed Neural Networks for Sparse Grid-Based Discontinuity Detectors

### v1

@@ 5.3 Costs and Potentialities
As observed in the experiments above, the proposed NN-based detectors have many potentialities. In particular, a well-trained NN-based $n$-dimensional detector can identify with high precision the discontinuity interfaces of generic discontinuous functions with domain in $\mathbb{R}^{n}$. We recall that the NN (better if a GINN) is trained only once on a synthetic dataset, showing interesting and very good generalization abilities. Moreover, such a kind of models have the advantage of being portable and versatile (see 5.2 below); i.e., they can be shared among different users and integrated also into algorithms different from the ones proposed in this paper. Furthermore, the experiments show that these detectors and algorithms are suitable for the discontinuity detection problem in dimensions higher than 3.
Nevertheless, some drawbacks still exist. The main one is the synthetic dataset creation, based on deterministic detectors; indeed, increasing the problem's dimensionality, the cost of this operation becomes particularly expensive. One solution is to parallelize the operation among more workers and gather the results; we adopted this approach for the creation of $\mathcal{D}_{4}$, reducing the cost from days to hours. However, we recall that the cost of the dataset creation is ``paid'' only once; hence, a price worth paying if it returns a good NN-based detector. Another drawback depends on the current implementation of Algorithms 1 and 2. Indeed, the current versions of these algorithms are not optimized yet with respect to the memory management and the check operation of the already-evaluated points; then, the procedure may slow down when the number of memorized points is large. However, this drawback can be overcome and considerably reduced in future work, optimizing the code implementation of Algorithms 1 and 2.
@@
The trained GINNs used in the experiments of Sections Section 5.1 and Section 5.2 are available at ADD LINK HERE.
@@ 6 Conclusion
In conclusion, our work presents a promising advancement in the field of discontinuity detection, particularly in domains of dimension higher than 3. Our novel approach utilizes Graph-Informed Neural Networks (GINNs) trained to detect troubled points within a sparse grid used for function evaluations, where the GINNs are based on the adjacency matrix of a graph constructed on the grid's structure. We have introduced a recursive algorithm for general sparse grid-based detectors, characterized by finite termination and convergence properties. Running this algorithm with a well-trained GINN as the detector results in a cost-effective and rapid discontinuity detection method suitable for functions with $n$-dimensional domains. Our experiments, conducted on functions with dimensions $n=2$ and $n=4$, demonstrate the efficiency and notable generalization abilities of the GINNs in detecting discontinuity interfaces, even for functions different from those used for building the training sets. The portable and versatile nature of the trained GINNs allows them to be shared among users and integrated into new algorithms. While our proposed NN-based detectors exhibit many potentials, some limitations persist. Creating synthetic datasets based on deterministic detectors can be costly, particularly as problem dimensionality increases. We have addressed this by parallelizing the dataset creation process. Moreover, we acknowledge that the current implementations of the detection algorithms may not be fully optimized in terms of memory management and check operations for already-evaluated points, potentially leading to slowdowns when the number of memorized points is significant. However, these drawbacks can be mitigated and reduced through future work focused on code optimization. In summary, our method offers a powerful tool for discontinuity detection, demonstrating its potential especially for dimensions higher than 3. Further research in dataset creation efficiency and algorithm optimization will permit to obtain even more efficient discontinuity detection techniques, expanding the applicability of this approach to a wider range of practical problems and, in particular, to even higher dimensions.
@@ Appendix A Practical Details for Building NN-based Detectors
In this appendix, we report the details of the procedure used for building the NN-based detectors. We start illustrating how to build a deterministic approximation of an exact detector $\Delta^{*}$ for discontinuous functions with discontinuities contained in the zero-level set of known continuous functions (Section A.1). Then, we introduce the random piece-wise continuous functions used for generating the synthetic data (Section A.2), the preprocessing function for NN inputs (Section A.3), and the architecture archetype used for the NN models (Section A.4).
@@ A.1 Zero-level Set Detection and Approximation of Exact Detectors
Let $\mathcal{C}$ denote the set of continuous functions $f:\Omega\rightarrow\mathbb{R}$, for any $\Omega\subseteq\mathbb{R}^{n}$. Then, for each fixed sparse grid $\mathcal{S}$, we can define a deterministic zero-level detector $Z^{(t+1)}:\mathcal{C}\times\mathcal{S}_{/\simeq}\rightarrow\{0,1\}^{N}$, $t\in\mathbb{N}$, $t\geq 2$, described by the following algorithm:
@@
    For each edge $\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}\in E^{\prime}$, $\boldsymbol{x}_{i}$ is detected as troubled point ($p_{i}=1$) if $s(\boldsymbol{x}_{i})=0$ or exist $\tau\in\{1,\ldots,\lceil t/2\rceil\}$ such that $s(\boldsymbol{x}_{i})\neq s(\boldsymbol{x}_{ij}^{(\tau)})$. Analogously, $\boldsymbol{x}_{j}$ is detected as troubled point if $s(f(\boldsymbol{x}_{j})=0$ or exist $\tau\in\{\lfloor t/2\rfloor,\ldots t-1\}$ such that $s(\boldsymbol{x}_{j})\neq s(\boldsymbol{x}_{ij}^{(\tau)})$.
@@
$\lim_{t\rightarrow+\infty}Z^{(t+1)}=Z^{*}$ (22)
@@
Let $e_{ij}\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}\in E^{\prime}$ be an edge intersecting $\zeta$ and let $\boldsymbol{x}_{i}^{\zeta}\in\zeta\cap e_{ij}$ be the nearest point to $\boldsymbol{x}_{i}$ such that $\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\zeta}}\|\leq\|{\boldsymbol{x}_{i}^{\zeta}-\boldsymbol{x}_{j}}\|$; i.e., by definition, $\boldsymbol{x}_{i}$ is a troubled point with respect to the zero-level set detection problem (generalization of 3.1).
@@ A.2 Synthetic Dataset Creation
      $\pi(\boldsymbol{x})=C\frac{p(\boldsymbol{\xi})}{\max_{\boldsymbol{y}\in[-1,1]^{n-1}}|p(\boldsymbol{y})|}-x_{\iota}\,,$
@@
      $p(\boldsymbol{\xi})=\sum_{\boldsymbol{h}\in\mathcal{I}_{\rm sum}(4)}\left(c_{\boldsymbol{h}}\prod_{i=1,i\neq\iota}^{n}P_{h_{i}}\left(\frac{x_{i}+1}{2}\right)\right)\,,$
@@
      Namely, $\pi$ is the polynomial $p$ rescaled to a maximum absolute value equal to C, minus $x_{\iota}$. See Figure 4 and Section 5 for 2D examples of such a kind of discontinuity interface.
@@ A.3 Preprocessing of Input Data
Let $\gamma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ denote the function representing the preprocessing operation of the input data. Naive but efficient preprocessing functions for our needs are the rescaling functions, because they permit to ``feed'' the NN with vectors having all the values into a chosen range $[\alpha,\beta]$. In particular, we use the following function:
@@ A.4 Architecture Archetypes
The architecture archetype we use is based on residual blocks [12] and exploits also the batch-normalization [13]. The number of units per layer is $N$ (i.e., the number of sparse grid nodes) while the depth is characterized by the diameter of the SGG; this depth characterization has been chosen mainly thinking to GINN models (see [5, Proposition 1]), but it is a good choice also for MLPs, since residual blocks permit to better exploit the model's depth (see [12]).
@@
  • One hidden GI/FC layer $L_{1}$ of $N$ units and activation function $\varphi$. If it is a GI layer, it is based on $A$ and it is characterized by $F\geq 1$ output features per node;
@@
    • One hidden GI/FC layer $L_{h+1}^{\prime}$ of $N$ units and activation function $\varphi$. If it is a GI layer, it is based on $A$ and it is characterized by $F\geq 1$ input and output features per node;
@@
    • One layer $\Sigma_{h+1}$ that returns the sum of the outputs of $L_{h+1}^{\prime\prime}$ and $\Sigma_{h}$ ($\Sigma_{1}\equiv L_{1}$ by convention), applying to it the activation function $\varphi$.
@@
Acknowledgements
This study was carried out within the FAIR - Future Artificial Intelligence Research and received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR) – MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3 – D.D. 1555 11/10/2022, PE00000013). This manuscript reflects only the authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them. S.P. acknowledges support from Italian MIUR PRIN project 20227K44ME, Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning (FaReX). The authors acknowledge support from INdAM-GNCS Group Projects 2023, CUP E53C22001930001.

### v4

@@ 5.3 Costs and Potentialities
As observed in the experiments above, the proposed NN-based detectors have good potential. In particular, a well-trained NN-based $n$-dimensional detector can identify with high precision the discontinuity interfaces of generic discontinuous functions with domain in $\mathbb{R}^{n}$. We recall that the NN (better if a GINN) is trained only once on a synthetic dataset, interestingly showing very good generalization abilities. Moreover, such a kind of models have the advantage of being portable and versatile (see Remark 5.2 below); i.e., they can be shared among different users and integrated also into algorithms different from the ones proposed in this paper. Furthermore, the experiments show that these detectors and algorithms are suitable for the discontinuity detection problem in dimensions higher than 3.
Nevertheless, some drawbacks still exist. The main one is the synthetic dataset creation, based on deterministic detectors; indeed, increasing the problem’s dimensionality, the cost of this operation becomes particularly expensive. One solution is to parallelize the operation; we adopted this approach for the creation of $\mathcal{D}_{4}$, reducing the cost from days to hours. However, we recall that the cost of the dataset creation is paid only once; hence, it is a price worth paying if it returns a good NN-based detector applicable to any discontinuity detection problem with the same domain dimension.
@@
The trained NNs used in the experiments of Section 5.1 and Section 5.2, an implementation of Algorithm 2, and practical toy examples are available at https://github.com/Fra0013To/NNbasedDiscDetectors. Users can modify the practical examples available in the repository to detect discontinuities in a custom target function.
@@ 6 Conclusion
Our work presents a promising advancement in the field of discontinuity detection, particularly in domains of dimension higher than 3. Our novel approach utilizes Graph-Instructed Neural Networks (GINNs) trained to detect troubled points within a sparse grid used for function evaluations, where the GINNs are based on the adjacency matrix of a graph constructed on the grid’s structure. We have introduced a recursive algorithm for general sparse grid-based detectors, characterized by finite termination and convergence properties. Running this algorithm with a well-trained GINN as the detector results in a cost-effective and rapid discontinuity detection method suitable for functions with $n$-dimensional domains. Our experiments, conducted on functions with dimensions $n=2$ and $n=4$, demonstrate the efficiency and notable generalization abilities of the GINNs in detecting discontinuity interfaces, even for functions different from those used for building the training sets. The portable and versatile nature of the trained GINNs allows them to be shared among users and integrated into new algorithms. While our proposed NN-based detectors exhibit a good potential, some limitations persist. Creating synthetic datasets based on deterministic detectors can be costly, particularly as problem dimensionality increases. We have addressed this by parallelizing the dataset creation process.
In summary, our method offers a powerful tool for discontinuity detection, demonstrating its potential, especially for dimensions higher than 3. Further research in dataset creation efficiency, optimal NN hyper-parameters search, and algorithm improvements will permit us to develop even more efficient software for discontinuity detection techniques, expanding the applicability of this approach to a wider range of practical problems and, in particular, to even higher dimensions.
Acknowledgements
The authors acknowledge that this study was carried out within the FAIR-Future Artificial Intelligence Research and received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR)–MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3—D.D. 1555 11/10/2022, PE00000013). This manuscript reflects only the authors’ views and opinions; neither the European Union nor the European Commission can be considered responsible for them. The authors acknowledge support from the Italian MIUR PRIN project 20227K44ME, Full and Reduced order modeling of coupled systems: focus on non-matching methods and automatic learning (FaReX).
Code Availability:
The code for loading the trained NN-based detectors illustrated in this paper, for an implementation of Algorithm 2, and for running practical toy examples is available at https://github.com/Fra0013To/NNbasedDiscDetectors. See also Remark 5.2.
@@ Appendix A Practical Details for Building NN-based Detectors
In this appendix, we report the details of the procedure used for building the NN-based detectors. We start illustrating how to build a deterministic approximation of an exact detector $\Delta^{*}$ for discontinuous functions with discontinuities contained in the zero-level set of known continuous functions (A.1). Then, we introduce the random piece-wise continuous functions used for generating the synthetic data (A.2), the preprocessing function for NN inputs (A.3), and the architecture archetype used for the NN models (A.4).
@@ A.1 Zero-level Set Detection and Approximation of Exact Detectors
Let $\mathcal{C}$ denote the set of continuous functions $f:\Omega\rightarrow\mathbb{R}$, for any $\Omega\subseteq\mathbb{R}^{n}$. Then, for each fixed sparse grid $\mathcal{S}$, we can define a deterministic zero-level detector $Z^{(t+1)}:\mathcal{C}\times\mathcal{S}_{/\simeq}\rightarrow\{0,1\}^{N}$, fixed $t\in\mathbb{N}$, $t\geq 2$, described by the following algorithm:
@@
    For each edge $\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}\in E^{\prime}$, $\boldsymbol{x}_{i}$ is detected as troubled point ($p_{i}=1$) if $s(\boldsymbol{x}_{i})=0$ or exist $\tau\in\{1,\ldots,\lceil t/2\rceil\}$ such that $s(\boldsymbol{x}_{i})\neq s(\boldsymbol{x}_{ij}^{(\tau)})$. Analogously, $\boldsymbol{x}_{j}$ is detected as troubled point if $s(\boldsymbol{x}_{j})=0$ or exist $\tau\in\{\lfloor t/2\rfloor,\ldots t-1\}$ such that $s(\boldsymbol{x}_{j})\neq s(\boldsymbol{x}_{ij}^{(\tau)})$.
@@
$\lim_{t\rightarrow+\infty}Z^{(t+1)}(\boldsymbol{x})=Z^{*}(\boldsymbol{x})\,,$ (22)
for each $\boldsymbol{x}\in\mathcal{S}^{\prime}$ and each $\mathcal{S}^{\prime}\simeq\mathcal{S}$.
@@
Let $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}\in E^{\prime}$ be an edge intersecting $\zeta$ and let $\boldsymbol{x}_{i}^{\zeta}\in\zeta\cap e_{ij}$ be the nearest point to $\boldsymbol{x}_{i}$ such that $\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\zeta}}\|\leq\|{\boldsymbol{x}_{i}^{\zeta}-\boldsymbol{x}_{j}}\|$; i.e., by definition, $\boldsymbol{x}_{i}$ is a troubled point with respect to the zero-level set detection problem (generalization of Definition 3.1).
@@ A.2 Synthetic Dataset Creation
      $\pi(\boldsymbol{x})=C\frac{\Pi(\boldsymbol{\xi})}{\max_{\boldsymbol{y}\in[-1,1]^{n-1}}|\Pi(\boldsymbol{y})|}-x_{\iota}\,,$
@@
      $\Pi(\boldsymbol{\xi})=\sum_{\boldsymbol{h}\in\mathcal{I}_{\rm sum}(4)}\left(c_{\boldsymbol{h}}\prod_{i=1,i\neq\iota}^{n}P_{h_{i}}\left(\frac{x_{i}+1}{2}\right)\right)\,,$
@@
      Namely, $\pi$ is the polynomial $\Pi$ rescaled to a maximum absolute value equal to C, minus $x_{\iota}$. See Figure 4 and Section 5 for 2D examples of such a kind of discontinuity interface.
@@ A.3 Preprocessing of Input Data
Let $\gamma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ denote the function representing the preprocessing operation of the input data. Naive but efficient preprocessing functions for our needs are the rescaling functions, because they permit to “feed” the NN with vectors having all the values into a chosen range $[\alpha,\beta]$. In particular, we use the following function:
@@ A.4 Architecture Archetypes
The architecture archetype we use is based on residual blocks [42] and exploits also the batch-normalization [44]. The number of units per layer is $N$ (i.e., the number of sparse grid nodes) while the depth is characterized by the diameter of the SGG; this depth characterization has been chosen mainly thinking to GINN models (see [26, Proposition 1]), but it is a good choice also for MLPs, since residual blocks permit to better exploit the model’s depth (see [42]).
@@
  • One hidden GI/FC layer $L_{1}$ of $N$ units and activation function $\psi$. If it is a GI layer, it is based on $A$ and it is characterized by $F\geq 1$ output features per node;
@@
    • One hidden GI/FC layer $L_{h+1}^{\prime}$ of $N$ units and activation function $\psi$. If it is a GI layer, it is based on $A$ and it is characterized by $F\geq 1$ input and output features per node;
@@
    • One layer $\Sigma_{h+1}$ that returns the sum of the outputs of $L_{h+1}^{\prime\prime}$ and $\Sigma_{h}$ ($\Sigma_{1}\equiv L_{1}$ by convention), applying to it the activation function $\psi$.